a man is seven times as old as his son. Ten years from now, he will be three times as old as his son. how old is he now?
Answer by Maths68(1474)

Let
Present age of Man = m
Present age of Son = s
A man is sevn times as old as his son.
m=7s........(1)
Ten year from now
m+10=s+10
He will be three times as old as his son
m+10=3(s+10)
m+10=3s+30
m=3s+30-10
m=3s+20........(2)
Subtitute the value of m from (1) to (2)
7s=3s+20
7s-3s=20
4s=20
4s/4=20/4
s=5
Put the value of s in (1)
m=7s
m=7(5)
m=35
Present age of Man = m = 35
Present age of Son = s = 5